### v1

@@ Appendix B Omitted Proofs of Section 4
Similarly, if $c\prec b\prec d\prec a$, then $b\prec f\prec a$ by the previous argument. Again, $f$ cannot be the right endpoint of $fg$ because then $fg$ would end below $cd$ or $ad$, and it cannot be the left endpoint of $fg$ by the same argument as before.
@@
Now suppose that $a\prec d\prec c\prec b$. By a similar argument as before, having $a\prec f\prec c$ leads to at least one forbidden configuration. Here, however, $f\prec a$ is possible because there is no edge ending at $a$ that could form a forbidden configuration with the lightest edge $df$ It follows that $a\prec e$ and $g\prec f$. It remains to give $e$ a precise position. If $a\prec e\prec b$, then $ae$ ends below one of the lighter edges $ad$, $cd$, or $bc$. However, if $b\prec e$, then the heavier edge $ab$ ends below $ae$.
@@
Finally suppose that $d\prec a\prec c\prec b$. We can argue again that $g\prec f\prec d\prec a\prec c\prec b$. It remains to place $e$. Clearly, $a\prec e$ as otherwise $ae$ ends below $cd$. Now, however, $ae$ ends below the lighter edges $cd$ or $cb$, or the heavier edge $ab$ ends below $ae$; a contradiction.
See 4.22
@@
Let $G=(V,E)$ be any fixed graph. Our task is to show that either $\operatorname{pqn}(G)=1$, or $G$ contains some $\mathcal{F}\in\{\mathcal{F}_{1},\ldots,\mathcal{F}_{8}\}$ as a minor. For this, we may assume that $G$ is connected. If $G$ is a tree, i.e., contains no cycles, then $\operatorname{pqn}(G)=1$ by Lemma 4.4 and we are done.
Next assume that $G$ contains exactly one cycle $C$. Then $G-E(C)$ is a forest $F$, and we consider each component of $F$ as rooted at the vertex of $C$. For a component $K$ of $F=G-E(C)$ we say that $K$ is
@@
First assume that some component $K$ of $F$ is complex. Let $w$ be the vertex in $K$ that lies on $C$. If $K$ is not a caterpillar, then $K$ contains a vertex $v$ (possibly $w$) with at least three neighbors of degree at least $2$ each. Contracting $C$ to a triangle and the $v$-$w$-path in $K$ to a single vertex results in a graph that contains $\mathcal{F}_{4}$. Hence, $G$ contains $\mathcal{F}_{4}$ as a minor. So assume that $K$ is a caterpillar but no underlying path of $K$ starts with $w$. Let $P$ be the path in $K$ on all non-leaf vertices of $K$. Then $w$ is neither an endpoint of $P$ nor adjacent to an endpoint of $P$. It follows that either $w$ is an interior vertex of $P$ or $w$ is adjacent to an interior vertex $w^{\prime}$ of $P$. Contracting $C$ again to a triangle, and contracting the edge $ww^{\prime}$ in the latter case, we again obtain a graph that contains $\mathcal{F}_{4}$. Hence again, $G$ contains $\mathcal{F}_{4}$ as a minor and we are left with the case that no component of $F$ is complex.
@@
Now if every component of $F$ is trivial or a rooted star, then $\operatorname{pqn}(G)=1$ by Lemma 4.10. So assume some component $K$ of $F$ is a rooted caterpillar. If all other components (different from $K$) are trivial, then $\operatorname{pqn}(G)=1$ by Lemma 4.12. If $F$ has besides $K$ two further non-trivial components $K^{\prime},K^{\prime\prime}\neq K$, then $G$ contains $\mathcal{F}_{5}$ as a minor. So assume now that there is exactly one further non-trivial component $K^{\prime}\neq K$. If the roots of $K$ and $K^{\prime}$ have distance at least $3$ in one direction along $C$, then $G$ contains $\mathcal{F}_{8}$ as a minor. So, if none of the above is the case, then either $C$ is a triangle and $\operatorname{pqn}(G)=1$ by Corollary 4.16, or $C$ is a quadrangle and the roots of $K$ and $K^{\prime}$ are opposite on $C$, and $\operatorname{pqn}(G)=1$ by Lemma 4.14.
@@
The case where $G$ contains at least two cycles remains. First assume that $G$ contains two edge-disjoint cycles $C_{1},C_{2}$. If $C_{1},C_{2}$ have at most one vertex in common, then $G$ contains $\mathcal{F}_{2}$ as a minor. If $C_{1},C_{2}$ have at least two vertices in common, we claim that $G$ contains $\mathcal{F}_{7}$ as a minor. One of $C_{1},C_{2}$, say $C_{2}$, has length at least $4$. Let $a$ and $c$ be two common vertices of $C_{1},C_{2}$ (consider Figure 4(g) for the naming scheme). Then $C_{1}$ can be contracted to the triangle $abc$, while $C_{2}$ and can be contracted to the 4-cycle $adce$. Omitting edge $ce$ yields $\mathcal{F}_{7}$.
@@
Finally, consider the case that any two cycles in $G$ share at least one edge. In fact, either $G$ contains $\mathcal{F}_{1}=K_{4}$ as a minor (and we are done) or there are two vertices $x,y$ in $G$ that are both contained in all cycles. Since there are no edge-disjoint cycles, there are at most three edge-disjoint $x$-$y$-paths in $G$. As $G$ has two cycles, there are exactly three such $x$-$y$-paths, say $P_{1},P_{2},P_{3}$. If $G$ is not just $P_{1}\cup P_{2}\cup P_{3}$, then $G$ contains $\mathcal{F}_{6}$ or $\mathcal{F}_{7}$ as a minor. If at least one of $P_{1},P_{2},P_{3}$ has at least three edges, then $G$ contains $\mathcal{F}_{3}$ as a minor. So, if none of the above is the case, then either $G=K_{2,3}$ or $G=K_{4}-e$ and we have $\operatorname{pqn}(G)=1$ by Lemmas 4.17 and 4.18.
Appendix C Omitted Content of Section 5
@@
Now, we present the full proof of Theorem 5.3.
@@
We define a vertex ordering $\sigma$ of a $2$-tree rooted at $e^{*}=v^{\prime}v^{\prime\prime}$ to be left-growing if no vertex (different from $v^{\prime}$, $v^{\prime\prime}$) appears to the right of both its parents.
@@
For each $k\geq 1$, any given edge $e^{*}$ of any given weight $w^{*}$, and any non-empty half-open interval $[x,y)\subset\mathbb{R}$, there exists a $2$-tree $H_{k}$ rooted at $e^{*}$ and an edge weighting $w_{k}$ of $H_{k}$ such that
@@
So for the remainder let us fix $k\geq 2$, and assume that the claim holds for $k-1$. Our task is to construct $\langle H_{k},w_{k}\rangle$ for the given root edge $e^{*}=v^{\prime}v^{\prime\prime}$, given weight $w^{*}$, and given interval $[x,y)$. By additive and multiplicative shifts of the weight function, we can assume for convenience that $[x,y)=[0,d+1)$ for $d=2k^{2}$. Starting with $e^{*}=v^{\prime}v^{\prime\prime}$ of weight $w_{k}(e^{*})=w^{*}$, we construct $\langle H_{k},w_{k}\rangle$ as follows. See Figure 5 for an illustration.
  • Stack $d$ vertices $a_{1},\ldots,a_{d}$ onto the root edge $e^{*}$, giving weight 00 to all $2d$ new edges.
@@
    • stack a vertex $b^{\prime}_{i}$ onto edge $a_{i}v^{\prime}$ giving weight 00 to both new edges,
    • let $a_{i}b^{\prime}_{i}$ be the root edge of $\langle H_{k-1},w_{k-1}\rangle$ with interval $[i,i+1)$, as well as the root of $\langle H_{k-1},w_{k-1}\rangle$ with interval $[d+1-i,d+1-i+1)$,
    • stack a vertex $b^{\prime\prime}_{i}$ onto edge $a_{i}v^{\prime\prime}$ giving weight 00 to both new edges, and
    • let $a_{i}b^{\prime\prime}_{i}$ be the root edge of $\langle H_{k-1},w_{k-1}\rangle$ with interval $[i,i+1)$, as well as the root of $\langle H_{k-1},w_{k-1}\rangle$ with interval $[d+1-i,d+1-i+1)$.
@@
By the Erdős–Szekeres theorem, there is a subset $I_{2}\subseteq I_{1}$ of at least $\sqrt{k^{2}}=k$ indices such that either $a_{i}\prec a_{j}$ for all $i,j\in I_{2}$ with $i<j$ or $a_{j}\prec a_{i}$ for all $i,j\in I_{2}$ with $i<j$. By symmetry, let us assume that $a_{j}\prec a_{i}$ for all $i,j\in I_{2}$ with $i<j$ [^8] We use the symmetry of the two copies of $H_{k-1}$ with root edge $b^{\prime}_{i}a_{i}$ here. In the assumed case, we focus on the copies of $H_{k-1}$ with intervals $[i,i+1)$. In the other case, we would consider the copies of $H_{k-1}$ with intervals $[d+1-i,d+1-i+1)$. We only need the interval weights to be increasing right-to-left. Note that for the largest index $\hat{i}=\max(I_{2})$ all vertices $a_{i}$ with $i\in I_{2}-\hat{i}$ lie below the edge $a_{\hat{i}}v^{\prime}$.
Now consider for each $i\in I_{2}-\hat{i}$ the copy of $2$-tree $H_{k-1}$ with root edge $b^{\prime}_{i}a_{i}$ and interval $[i,i+1)$. By induction, for each $i\in I_{2}-\hat{i}$, there is a $(k-1)$-inversion whose right end is $a_{i}$ and all of whose edge-weights lie in the interval $[i,i+1)$.
Now, if for at least one such $i$, the left end of the corresponding $(k-1)$-inversion $(e_{1},\ldots,e_{k-1})$ lies right of $a_{\hat{i}}$, then $(a_{\hat{i}}v^{\prime},e_{1},\ldots,e_{k-1})$ is the desired $k$-inversion, and we are done. Otherwise, each of the first edges from the $k-1$ corresponding $(k-1)$-inversions has its left endpoint to the left of $a_{\hat{i}}$. But then, taking edge $a_{\hat{i}}v^{\prime}$ and these $k-1$ first edges in order of increasing weights again gives a desired $k$-inversion.
To finish the proof of the theorem, it is now enough to take $\langle H_{p},w_{p}\rangle$ from 3 with any interval $[x,y)$, and then apply 1 to obtain the desired edge-weighted $2$-tree $\langle G_{p},w_{p}\rangle$ with $\operatorname{pqn}(G_{p},w_{p})\geq p$.
Appendix D Omitted Proofs of Section 6
It remains to prove 2, which is part of the proof of Theorem 6.1.

### v3

@@ Appendix B Omitted Proofs of Section˜4
Second, suppose that $c\prec b\prec d\prec a$. Then again no vertex from $x^{*}$ on the path between $b$ and $c$ lies to the right of $d$. By swapping the roles of $b$ and $c$, we can use the same argument as before to obtain $b\prec f\prec a$. Again, $f$ cannot be the right endpoint of $fg$ because then $fg$ would end below the lighter edge $cd$ or below a lighter edge on the path from $d$ to $a$, and it cannot be the left endpoint of $fg$ by the same argument as before.
@@
Third, suppose that $a\prec d\prec c\prec b$. We next locate $f$. Since $c$ is the right endpoint of the edge $cd$ with weight 3, it follows that $f\prec c$ as the edge $df$ has weight 1. Suppose that $a\prec f\prec c$. As before, $f$ cannot be the right endpoint of $fg$ because then the edge $fg$ would end below a lighter edge on the path from $a$ to $d$ or below the lighter edge $cd$. Also, $f$ cannot be the left endpoint of $fg$ because then, if $f\prec g\prec b$, the edge $fg$ would end below the lighter edge $cd$ or below a lighter edge on the path from $c$ to $b$, or, if $f\prec b\prec g$, the heavier edge $ab$ would end below $fg$. So, suppose that $f\prec a$. It follows that $a\prec e$ and $g\prec f$. It remains to give $e$ a precise position. If $a\prec e\prec b$, then $ae$ ends below one of the lighter edges on the path from $a$ to $d$, or below $cd$, or below one of the lighter edges on the path from $c$ to $d$. However, if $b\prec e$, then the heavier edge $ab$ ends below $ae$.
@@
Finally, suppose that $d\prec a\prec c\prec b$. Observe that we can use similar arguments as in the previous case to show that $g\prec f\prec d\prec a\prec c\prec b$. It remains to place $e$. Clearly, $a\prec e$ as otherwise $ae$ ends below $cd$. Now, however, $ae$ ends below the lighter edge $cd$ or a lighter edge on the path from $c$ to $b$, or the heavier edge $ab$ ends below $ae$; a contradiction.
$\mathcal{F}_{5}$, $\mathcal{F}_{6}$
: We first consider the case $G\in\{\mathcal{F}_{5},\mathcal{F}_{6}\}$. By Lemma˜4.15 and symmetry, we may assume that $a,b\prec c$. Since each edge not in $\triangle abc$ has a weight strictly between $w(ab)=w(ac)$ and $w(bc)$, no such edge can start weakly left of $\min\{a,b\}$ and end strictly right of $\min\{a,b\}$. In particular, it follows that $d\prec a$ and, since also $w(ed)<w(ad)$, that $e\prec d$. In summary, we have $e\prec d\prec a\prec c$ and $b\prec c$.
Now for $\mathcal{F}_{5}$ it follows that the right endpoint of edge $fg$ must be left of $e$. But then edge $ed$ nests below $af$; a contradiction, as $w(ed)>w(af)$.
For $\mathcal{F}_{6}$, first note that if $a\prec b$, then either edge $bf$ ends below $ac$ (if $f\prec c$) or edge $bc$ ends below $bf$ (if $c\prec f$); either is not allowed. So we have $b\prec a$, and hence $f\prec a$, as argued in the beginning. But now either edge $bf$ ends below $ad$ (if $d\prec b$) or edge $ed$ ends below $ab$ (if $b\prec d$); either is not allowed.
Consider now the case where $G$ contains an $\mathcal{F}_{5}$ minor, i.e., $G$ contains a subgraph isomorphic to a subdivision of $\mathcal{F}_{5}$. That is, $G$ consists of a cycle $C$ that contains a vertex $a$ and a non-adjacent edge $bc$, a simple path $P_{a}$ from $a$ to a vertex $a^{\prime}$ that contains no edge of $C$ and two paths $P_{a^{\prime}e}=a^{\prime}de$ and $P_{a^{\prime}g}=a^{\prime}fg$ distinct from $C$ and $P_{a}$. We construct the weight function $w$ similar as before, namely, we set $w(bc)=6$, otherwise $w(e)=0$ for each edge $e$ on $C$, $w(e_{a})=1$ for each edge $e_{a}$ on $P_{a}$, and $w(a^{\prime}d)=4$, $w(de)=3$, $w(a^{\prime}f)=2$, $w(fg)=5$. Once more, by Lemma˜4.15 and symmetry, we may assume that $a,b\prec c$ as $bc$ is the heaviest edge on $C$.
We next claim that $a^{\prime}\prec a$. First observe that $^{\prime}a\prec c$ as otherwise edge $bc$ ends at $c$ but there is at least one edge $e_{a}$ of the path $P_{a}$ being co-active with $bc$ that ends after $c$ which is a contradiction to the fact that $w(e_{a})<w(bc)$. Assume now for a contradiction that $a\prec a^{\prime}\prec c$ and consider the simple path $P_{ac}$ between $a$ and $c$ along $C$ whose edges all have weight 0. Since $P_{a}$ and $P_{ac}$ both start at $a$ and $a^{\prime}precc$ there must be an edge $e_{a}$ of $P_{a}$ and an edge $e_{ac}$ of $P_{ac}$ that are co-active and such that $e_{a}$ ends before $e_{a}c$. This is however a contradiction to the fact that $w(e_{ac})<w(e_{a})$.
Hence, we can assume from now on that $a^{\prime}\prec a$. As for the case of $\mathcal{F}_{5}$, we can now observe that all of edges $a^{\prime}d$, $de$, $a^{\prime}f$ and $fg$ have higher weights then any of the already drawn edges except for $bc$. Thus, none of these edges can start weakly left of $\min\{a^{\prime},b\}$ and end strictly right of $\min\{a^{\prime},b\}$. Hence, $e,d,f,g\prec a^{\prime}$ and $e\prec d$. As for $\mathcal{F}_{5}$, the right endpoint of $fg$ must be left of $e$ which leads to $ed$ nesting $a^{\prime}f$; a contradiction.
Finally, consider the case where $G$ contains an $\mathcal{F}_{6}$ minor, i.e., it contains a subdivision of $\mathcal{F}_{6}$ as a minor. In this scenario, we observe that one of the following two scenarios happens: If we only subdivide any of the edges $ad$, dede, bfbf or $cg$, we obtain a graph that again contains $\mathcal{F}_{6}$ as a subgraph. Otherwise, we subdivide the triangle $\triangle abc$. However, subdividing the triangle $\triangle abc$ once at a single edge provides a supergraph of $\mathcal{F}_{4}$. Thus, any subdivision leads either to an $\mathcal{F}_{6}$ subgraph or to an $\mathcal{F}_{4}$ minor which concludes the case.
$\mathcal{F}_{7}$, $\mathcal{F}_{8}$
: We first consider the case where $G\in\{\mathcal{F}_{7},\mathcal{F}_{8}\}$. By Lemma˜4.15, $b$ or $c$ is the last vertex of the 4-cycle abcdabcd and $a$ or $c$ is the last vertex of $\triangle abc$. Hence, $a,b,d\prec c$. Moreover, $a\prec d$ since otherwise, $ad$ would end below the lighter edge $cd$. We know that $e\prec c$ because otherwise $ac$ would end below the lighter edge dede (aeae, respectively).
For $\mathcal{F}_{7}$, further $e\prec d$ because if $d\prec e\prec c$, then $de$ would end below the lighter edge $cd$. It remains to place $b$ (we already know $b\prec c$). It cannot be $b\prec d$ because then $de$ would end below the lighter edge $bc$. Hence, we have $a\prec d\prec b\prec c$. Now, however, $ab$ ends below the lighter edge $cd$; a contradiction.
For $\mathcal{F}_{8}$, a similar argument applies. We have $e\prec a$ because if $a\prec e\prec c$, then $ae$ would end below one of the lighter edges $ad$ or $cd$. If $b\prec a\prec d\prec c$, then $ae$ would end below the lighter edge $bc$. If $a\prec b\prec d\prec c$, then $ab$ would end below the lighter edge $ad$. Hence, we have $a\prec d\prec b\prec c$. Now, however, $ab$ ends below the lighter edge $cd$; a contradiction.
Finally, consider the case where $G$ has a proper $\mathcal{F}_{7}$ or $\mathcal{F}_{8}$ minor, i.e., it contains a subgraph isomorphic to subdivision of $\mathcal{F}_{7}$ or $\mathcal{F}_{8}$, respectively. If the subdivision is obtained only by subdividing edge $de$ or $ae$, the obtained graph contains $\mathcal{F}_{7}$ or $\mathcal{F}_{8}$, respectively. Otherwise, if any of the edges $ab$, bcbc, cdcd or $da$ are subdivided, the graph contains a subdivision of $\mathcal{F}_{1}$.
Hence, we may assume that the only subdivided edge is $ac$. Moreover, we can assume that $ac$ is subdivided exactly once as otherwise we again yield a graph that contains a $\mathcal{F}_{1}$ minor. Let $f$ denote the subdivision vertex of $ac$ and let $w(af)$ and $w(cf)=2$.
By Lemma˜4.15, $b$ or $c$ is the last vertex of the $4$-cycle abcdabcd and $a$ or $f$ is the last vertex of the $4$-cycle abcfabcf. It follows, that $a,b,c,d\prec f$. Next, we observe that $b\prec c$ as otherwise the edges $cb$ and $cf$ start at $c$, i.e., they are co-active, but $bc$ ends before $cf$ despite $w(bc)>w(cf)$. Hence, $c$ is the last vertex of the $4$-cycle abcdabcd and as previously we obtain $a,b,d\prec c$. Moreover, as before $a\prec d$ since otherwise, $ad$ would end below the lighter edge co-active edge $cd$.
Up to this point, we established that $a\prec d\prec c\prec f$. We now place vertex $b$. If $a\prec b\prec c$, edge $ab$ would be co-active with and end before either $ad$ or $dc$ despite the fact that $w(ab)>w(ad)>w(dc)$. On the other hand, if $c\prec b\prec f$, edges $bc$ and $cf$ are co-active and the heavier edge $bc$ ends before the lighter edge $cf$. Finally, if $f\prec b$, the heavier edge $af$ ends before the lighter co-active edge $bc$. Hence, we conclude that $b\prec a\prec d\prec c\prec f$.
Now consider the placement of $e$. If $e\prec c$, the edge $de$ or $ae$ (in the subdivision of $\mathcal{F}_{7}$ or $\mathcal{F}_{8}$, respectively) is co-active with $bc$ but ends before it even though it is heaver than $bc$. If $c\prec e\prec f$, edge $de$ or $ae$ is co-active with the lighter edge $cf$ but ends before it. Finally, if $f\prec e$, edge $de$ or $ae$ is co-active with the heavier edge $af$ but ends after it. Thus, there is no valid placement for $e$ and $pqn(G)>1$.
See 4.18
@@
Let $G=(V,E)$ be any fixed graph. Our task is to show that either $\operatorname{pqn}(G)=1$, or $G$ contains some $\mathcal{F}\in\{\mathcal{F}_{1},\ldots,\mathcal{F}_{8}\}$ as a minor. For this, we may assume that $G$ is connected. If $G$ is a tree, i.e., contains no cycles, then $\operatorname{pqn}(G)=1$ by Lemma˜4.1 and we are done.
Next assume that $G$ contains exactly one cycle $C$. Then GE(C)G-E(C) is a forest $F$, and we consider each component of $F$ as rooted at the vertex of $C$. For a component $K$ of F=GE(C)F=G-E(C) we say that $K$ is
@@
First assume that some component $K$ of $F$ is complex. Let $w$ be the vertex in $K$ that lies on $C$. If $K$ is not a caterpillar, then $K$ contains a vertex $v$ (possibly $w$) with at least three neighbors of degree at least $2$ each. Contracting $C$ to a triangle and the $v$-$w$-path in $K$ to a single vertex results in a graph that contains $\mathcal{F}_{5}$. Hence, $G$ contains $\mathcal{F}_{5}$ as a minor. So assume that $K$ is a caterpillar but no underlying path of $K$ starts with $w$. Let $P$ be the path in $K$ on all non-leaf vertices of $K$. Then $w$ is neither an endpoint of $P$ nor adjacent to an endpoint of $P$. It follows that either $w$ is an interior vertex of $P$ or $w$ is adjacent to an interior vertex $w^{\prime}$ of $P$. Contracting $C$ again to a triangle, and contracting the edge $ww^{\prime}$ in the latter case, we again obtain a graph that contains $\mathcal{F}_{5}$. Hence again, $G$ contains $\mathcal{F}_{5}$ as a minor and we are left with the case that no component of $F$ is complex.
@@
Now if every component of $F$ is trivial or a rooted star, then $\operatorname{pqn}(G)=1$ by Lemma˜4.7. So assume some component $K$ of $F$ is a rooted caterpillar. If all other components (different from $K$) are trivial, then $\operatorname{pqn}(G)=1$ by Lemma˜4.9. If $F$ has besides $K$ two further non-trivial components $K^{\prime},K^{\prime\prime}\neq K$, then $G$ contains $\mathcal{F}_{6}$ as a minor. So assume now that there is exactly one further non-trivial component $K^{\prime}\neq K$. If the roots of $K$ and $K^{\prime}$ have distance at least $3$ in one direction along $C$, then $G$ contains $\mathcal{F}_{4}$ as a minor. So, if none of the above is the case, then either $C$ is a triangle and $\operatorname{pqn}(G)=1$ by Lemma˜4.11, or $C$ is a quadrangle and the roots of $K$ and $K^{\prime}$ are opposite on $C$, and $\operatorname{pqn}(G)=1$ by Lemma˜4.11.
@@
The case where $G$ contains at least two cycles remains. First assume that $G$ contains two edge-disjoint cycles $C_{1},C_{2}$. If $C_{1},C_{2}$ have at most one vertex in common, then $G$ contains $\mathcal{F}_{2}$ as a minor. If $C_{1},C_{2}$ have at least two vertices in common, we claim that $G$ contains $\mathcal{F}_{8}$ as a minor. One of $C_{1},C_{2}$, say $C_{2}$, has length at least $4$. Let $a$ and $c$ be two common vertices of $C_{1},C_{2}$ (consider Figure˜4(g) for the naming scheme). Then $C_{1}$ can be contracted to the triangle abcabc, while $C_{2}$ and can be contracted to the 4-cycle adceadce. Omitting edge $ce$ yields $\mathcal{F}_{8}$.
@@
Finally, consider the case that any two cycles in $G$ share at least one edge. In fact, either $G$ contains $\mathcal{F}_{3}=K_{4}$ as a minor (and we are done) or there are two vertices $x,y$ in $G$ that are both contained in all cycles. Since there are no edge-disjoint cycles, there are at most three edge-disjoint $x$-$y$-paths in $G$. As $G$ has two cycles, there are exactly three such xxyy-paths, say $P_{1},P_{2},P_{3}$. If $G$ is not just $P_{1}\cup P_{2}\cup P_{3}$, then $G$ contains $\mathcal{F}_{7}$ or $\mathcal{F}_{8}$ as a minor. If at least one of $P_{1},P_{2},P_{3}$ has at least three edges, then $G$ contains $\mathcal{F}_{1}$ as a minor. So, if none of the above is the case, then either $G=K_{2,3}$ or $G=K_{4}-e$ and we have $\operatorname{pqn}(G)=1$ by Lemmas˜4.13 and 4.14.
Appendix C Omitted Content of Section˜5
@@
Now, we present the full proof of Theorem˜5.3.
@@
We define a vertex ordering $\sigma$ of a $2$-tree rooted at $e^{*}=v^{\prime}v^{\prime\prime}$ to be left-growing if no vertex (different from $^{\prime}, v′′v^{\prime\prime}$) appears to the right of both its parents.
@@
For each $k\geq 1$, any given edge $e^{*}$ of any given weight $w^{*}$, and any non-empty half-open interval $x,y)\subset\mathbb{R}$, there exists a $2$-tree $H_{k}$ rooted at $e^{*}$ and an edge weighting $w_{k}$ of $H_{k}$ such that
@@
So for the remainder let us fix $k\geq 2$, and assume that the claim holds for k1k-1. Our task is to construct $\langle H_{k},w_{k}\rangle$ for the given root edge $e^{*}=v^{\prime}v^{\prime\prime}$, given weight $w^{*}$, and given interval $[x,y)$. By additive and multiplicative shifts of the weight function, we can assume for convenience that $[x,y)=[0,d+1)$ for $d=2k^{2}$. Starting with $e^{*}=v^{\prime}v^{\prime\prime}$ of weight $w_{k}(e^{*})=w^{*}$, we construct $\langle H_{k},w_{k}\rangle$ as follows. See Figure˜5 for an illustration.
  • Stack $d$ vertices $a_{1},\ldots,a_{d}$ onto the root edge $e^{*}$, giving weight 0 to all $2d$ new edges.
@@
    • stack a vertex $b^{\prime}_{i}$ onto edge $a_{i}v^{\prime}$ giving weight 0 to both new edges,
    • let $a_{i}b^{\prime}_{i}$ be the root edge of $\langle H_{k-1},w_{k-1}\rangle$ with interval $[i,i+1)$, as well as the root of $\langle H_{k-1},w_{k-1}\rangle$ with interval [d+1i,d+1i+1)[d+1-i,d+1-i+1),
    • stack a vertex $b^{\prime\prime}_{i}$ onto edge $a_{i}v^{\prime\prime}$ giving weight 0 to both new edges, and
    • let $a_{i}b^{\prime\prime}_{i}$ be the root edge of $\langle H_{k-1},w_{k-1}\rangle$ with interval $[i,i+1)$, as well as the root of $\langle H_{k-1},w_{k-1}\rangle$ with interval [d+1i,d+1i+1)[d+1-i,d+1-i+1).
@@
By the Erdős–Szekeres theorem, there is a subset $I_{2}\subseteq I_{1}$ of at least $\sqrt{k^{2}}=k$ indices such that either $a_{i}\prec a_{j}$ for all $i,j\in I_{2}$ with $i<j$ or $a_{j}\prec a_{i}$ for all $i,j\in I_{2}$ with $i<j$. By symmetry, let us assume that $a_{j}\prec a_{i}$ for all $i,j\in I_{2}$ with i<ji<j. [^7] We use the symmetry of the two copies of $H_{k-1}$ with root edge $b^{\prime}_{i}a_{i}$ here. In the assumed case, we focus on the copies of $H_{k-1}$ with intervals $[i,i+1)$. In the other case, we would consider the copies of $H_{k-1}$ with intervals [d+1i,d+1i+1)[d+1-i,d+1-i+1). We only need the interval weights to be increasing right-to-left. Note that for the largest index $\hat{i}=\max(I_{2})$ all vertices $a_{i}$ with $i\in I_{2}-\hat{i}$ lie below the edge $a_{\hat{i}}v^{\prime}$.
Now consider for each $i\in I_{2}-\hat{i}$ the copy of $2$-tree $H_{k-1}$ with root edge $b^{\prime}_{i}a_{i}$ and interval $[i,i+1)$. By induction, for each $i\in I_{2}-\hat{i}$, there is a (k1)(k-1)-inversion whose right end is $a_{i}$ and all of whose edge-weights lie in the interval $[i,i+1)$.
Now, if for at least one such $i$, the left end of the corresponding (k1)(k-1)-inversion $(e_{1},\ldots,e_{k-1})$ lies right of $a_{\hat{i}}$, then $(a_{\hat{i}}v^{\prime},e_{1},\ldots,e_{k-1})$ is the desired $k$-inversion, and we are done. Otherwise, each of the first edges from the k1k-1 corresponding (k1)(k-1)-inversions has its left endpoint to the left of $a_{\hat{i}}$. But then, taking edge $a_{\hat{i}}v^{\prime}$ and these k1k-1 first edges in order of increasing weights again gives a desired $k$-inversion.
To finish the proof of the theorem, it is now enough to take $\langle H_{p},w_{p}\rangle$ from ˜3 with any interval $[x,y)$, and then apply ˜1 to obtain the desired edge-weighted $2$-tree $\langle G_{p},w_{p}\rangle$ with $\operatorname{pqn}(G_{p},w_{p})\geq p$.
Appendix D Omitted Proofs of Section˜6
It remains to prove ˜2, which is part of the proof of Theorem˜6.1.